OpenStudy (torii897):

PLEASE I NEED HELP iv tried so many time 7 + (10 − 4)2 ÷ 4 × 1 over 2 to the power of 3? A:8 B:8 and 1 over 6 C:8 and 1 over 8 D:79

Mehek (mehek14):

\(7+(10-4)^2÷4*\dfrac{1}{2}^3\)

Mehek (mehek14):

10-4 = 6 \(6^2=36\\ 7 + 36 ÷ 4 * \dfrac{1}{2}^3\\7+36÷4*\dfrac{1}{8}\)

Mehek (mehek14):

\(36÷4=9\)

Mehek (mehek14):

\(\dfrac{9}{1}*\dfrac{1}{8}=\dfrac{9}{8}=1\dfrac{1}{8}\)

Mehek (mehek14):

\(7+1\dfrac{1}{8} = 8 \dfrac{1}{8}\)

Mehek (mehek14):

so C

Mehek (mehek14):

@torii897

OpenStudy (torii897):

thank you sooo much!!!!!!

Mehek (mehek14):

yw

OpenStudy (torii897):

can u help me with a few more?